# Phase transitions for unique codings of fat Sierpinski gasket with multiple digits

### v1

@@ Abstract.
Given an integer $M\geq 1$ and $\beta\in(1,M+1)$, let $S_{\beta,M}$ be the fat Sierpinski gasket in $\mathbb{R}^{2}$ generated by the iterated function system $\left\{f_{d}(x)=\frac{x+d}{\beta}:d\in\Omega_{M}\right\}$, where $\Omega_{M}=\{(i,j)\in\mathbb{Z}_{\geq 0}^{2}:i+j\leq M\}$. Then each $x\in S_{\beta,M}$ may represent as a series $x=\sum_{i=1}^{\infty}\frac{d_{i}}{\beta^{i}}=:\Pi_{\beta}((d_{i}))$, and the infinite sequence $(d_{i})\in\Omega_{M}^{\mathbb{N}}$ is called a coding of $x$. Since $\beta<M+1$, a point in $S_{\beta,M}$ may have multiple codings. Let $U_{\beta,M}$ be the set of $x\in S_{\beta,M}$ having a unique coding, that is
@@
When $M=1$, Kong and Li [2020, Nonlinearity] described two critical bases for the phase transitions of the intrinsic univoque set $\widetilde{U}_{\beta,1}$, which is a subset of $U_{\beta,1}$. In this paper we consider $M\geq 2$, and characterize the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ for the phase transitions of $U_{\beta,M}$: (i) if $\beta\in(1,\beta_{G}(M)]$, then $U_{\beta,M}$ is finite; (ii) if $\beta\in(\beta_{G}(M),\beta_{c}(M))$ then $U_{\beta,M}$ is countably infinite; (iii) if $\beta=\beta_{c}(M)$ then $U_{\beta,M}$ is uncountable and has zero Hausdorff dimension; (iv) if $\beta>\beta_{c}(M)$ then $U_{\beta,M}$ has positive Hausdorff dimension. Our results can also be applied to the intrinsic univoque set $\widetilde{U}_{\beta,M}$. Moreover, we show that the first critical base $\beta_{G}(M)$ is a perron number, while the second critical base $\beta_{c}(M)$ is a transcendental number.
@@
Fat Sierpinski gasket, unique coding, phase transition, critical bases, Thue-Morse sequence.
@@ 1. Introduction
It is well-known that $S_{\beta,M}$ is the unique nonempty compact set satisfying the equation $S_{\beta,M}=\bigcup_{d\in\Omega}f_{\beta,d}(S_{\beta,M})$ (cf. [18]). To study $S_{\beta,M}$ it is useful to introduce the coding map
@@
which is always continuous and surjective. If $\beta>M+1$, the IFS $\mathscr{F}_{\beta,M}=\left\{f_{\beta,d}:d\in\Omega_{M}\right\}$ satisfies the strong separation condition, and then $\Pi_{\beta}$ is bijective (cf. [15]). If $\beta=M+1$, the IFS $\mathscr{F}_{\beta,M}$ fails the strong separation condition but still admits the open set condition, and thus $\Pi_{\beta}$ is nearly bijective up to a countable set. However, for $\beta<M+1$ the IFS $\mathscr{F}_{\beta,M}$ fails the open set condition and it always has overlaps. In particular, if $\beta\in(1,\frac{M+2}{2}]$ the attractor $S_{\beta,M}$ is an isosceles triangle with vertices $(0,0),(\frac{M}{\beta-1},0)$ and $(0,\frac{M}{\beta-1})$ (see Proposition 2.5). In this case, for Lebesgue almost every $x\in S_{\beta,M}$ the preimage $\Pi_{\beta}^{-1}(x)$ is uncountable (cf. [28]). When $\beta\in(\frac{M+2}{2},M+1)$ the structure of $S_{\beta,M}$ is getting more complicated. Broomhead et al [5] showed that for $M=1$ and $\beta\leq\beta_{*}$ the set $S_{\beta,M}$ has non-empty interior, where $\beta_{*}\approx 1.54369$ is a root of $x^{3}-2x^{2}+2x-2=0$.
@@
$x$$y$$\alpha_{00}$$\alpha_{10}$$\alpha_{20}$$\alpha_{30}$$\alpha_{40}$$\alpha_{01}$$\alpha_{02}$$\alpha_{03}$$\alpha_{04}$$\alpha_{11}$$\alpha_{12}$$\alpha_{13}$$\alpha_{21}$$\alpha_{31}$$\alpha_{22}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{4}{\beta}$$\frac{4}{\beta-1}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{4}{\beta}$$\frac{4}{\beta-1}$
@@
(1.2) $\widetilde{U}_{\beta,M}:=\left\{\Pi_{\beta}((d_{i}))\in S_{\beta,M}:\Pi_{\beta}(d_{n+1}d_{n+2}\ldots)\notin O_{\beta,M}\leavevmode\nobreak\ \forall n\geq 0\right\},$
where $\Pi_{\beta}$ is defined in (1.1). Note that $\widetilde{U}_{\beta,M}$ is the survivor set of some open dynamical systems (cf. [27]). Indeed, let $T_{\beta,M}$ be the expanding map from the isosceles triangle $\Delta_{\beta,M}$ to itself such that $T_{\beta,M}(x)=f_{\beta,d}^{-1}(x)$ if $x\in f_{\beta,d}(\Delta_{\beta,M})$ for any $d\in{\Omega}_{M}$. Then $S_{\beta,M}$ is invariant under $T_{\beta,M}$, and the intrinsic univoque set $\widetilde{U}_{\beta,M}$ is the survivor set consisting of all $x\in\Delta_{\beta,M}$ whose orbit under $T_{\beta,M}$ never enters the hole $O_{\beta,M}$, i.e., $\widetilde{U}_{\beta,M}=\left\{x\in\Delta_{\beta,M}:T_{\beta,M}^{n}(x)\notin O_{\beta,M}\leavevmode\nobreak\ \forall n\geq 0\right\}$. For more information on open dynamical systems we refer to the papers [30, 12, 13, 11, 16, 7, 6, 20, 1] and the references therein.
@@
where #$\#A$ denotes the cardinality of a set $A$. Note that each point in $\widetilde{U}_{\beta,M}$ has a unique coding under $\Pi_{\beta}$. Then $\widetilde{U}_{\beta,M}\subset U_{\beta,M}$ for any $\beta\in(1,M+1)$.
@@
Denote by :=$\mathbb{N}:=\left\{1,2,3,\ldots\right\}$ and $\mathbb{N}_{0}:=\mathbb{N}\cup\left\{0\right\}$. Let $(\tau_{i})_{i=0}^{\infty}=01101001\ldots\in\left\{0,1\right\}^{\mathbb{N}}$ be the classical Thue-Morse sequence defined by (cf. [2])
@@
where $(\tau_{i})_{i=0}^{\infty}=01101001\ldots$ is the classical Thue-Morese sequence. Similarly, if $M=3N+2$ with $N\in\mathbb{N}_{0}$, then $\beta_{c}(3N+2)$ is the unique root in $(N+1,N+2)$ of the equation
@@
$[\frac{M}{3}]+1\leq\beta_{G}(M)<\beta_{c}(M)<[\frac{M}{3}]+2,$
@@
We will show that each $\beta_{G}(M)$ is a Perron number, and each $\beta_{c}(M)$ is a transcendental number. When $M=1$ the following phase transition phenomenon for $\widetilde{U}_{\beta,1}$ was established by Sidorov [28, Theorem 4.1] and Kong and Li [23, Theorem 2].
@@
    If $\beta\in(1,\beta_{G}(1)]$, then #$\#\widetilde{U}_{\beta,1}=\#U_{\beta,1}=3$;
@@
In this paper we extend Theorem 1.3 to $M\geq 2$, and show that (i) $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$ are finite if and only if $\beta\leq\beta_{G}(M)$; (ii) $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$ have positive Hausdorff dimension if and only if $\beta>\beta_{c}(M)$. The new challenge is that for $M=3N+1$ with $N\geq 1$ the digit $N-1$ might occur in $d_{n}^{1}$ or $d_{n}^{2}$ for a sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$, while for $M=1$ this can never happen. Fortunately, we are able to show that this digit $N-1$ can occur at most 3 times in any sequence of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ (see Lemma 4.5). Furthermore, completely different from the case for $M=1$, we need more effort to find the critical base $\beta_{c}(M)$ for $M=3N+3$. In other words, we need to construct a completely new type of sequences to represent $\beta_{c}(M)$, see more in Section 6.
@@
    If $\beta\in(1,\beta_{G}(M)]$, then #$\#U_{\beta,M}=\#\widetilde{U}_{\beta,M}=3$;
@@
    If $\beta>\beta_{c}(M)$, then both $U_{\beta,M}$ and $\widetilde{U}_{\beta,M}$ have positive Hausdorff dimensions.
@@
The rest of the paper is organized as follows. In the next section we give the characterization of $\widetilde{U}_{\beta,M}$. In Section 3 we show that the generalized golden ratio $\beta_{G}(M)$ defined in Definition 1.1 is the first critical base for $\widetilde{U}_{\beta,M}$ that $\widetilde{U}_{\beta,M}$ is finite if and only if $\beta\leq\beta_{G}(M)$. Furthermore, we show that the generalized Komornik-Loreti constant $\beta_{c}(M)$ in Definition 1.2 is the second critical base for $\widetilde{U}_{\beta,M}$. This will be proved in Sections 46 for $M=3N+1,M=3N+2$ and $M=3N+3$, respectively. We point out that the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ are also applied to $U_{\beta,M}$ by Proposition 2.5.
@@ 2. Characterization of $\widetilde{U}_{\beta,M}$
This representation plays an important role in our proof of Theorem 1.4; and it will be clear after we proving Proposition 2.3. Note by (2.1) that each row in the above presentation is a length $n$ word or an infinite sequence over the alphabet $\left\{0,1,\ldots,M\right\}$.
@@
Note that the two induced sequences $(d_{i}^{1})$ and $(d_{i}^{2})$ both have digits in $\left\{0,1,\ldots,M\right\}$. Then we also need to introduce the one-dimensional symbolic dynamics ($(\left\{0,1,\ldots,M\right\}^{\mathbb{N}},\sigma)$, where $\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ is the set of all infinite sequences with each digit from $\left\{0,1,\ldots,M\right\}$, and $\sigma$ is the left shift map. For $n\in\mathbb{N}$ let $\left\{0,1,\ldots,M\right\}^{n}$ be the set of all length $n$ words with each digit from $\left\{0,1,\ldots,M\right\}$. Set $\left\{0,1,\ldots,M\right\}^{*}:=\bigcup_{n=1}^{\infty}\left\{0,1,\ldots,M\right\}^{n}$. For a word $w=w_{1}\ldots w_{n}\in\left\{0,1,\ldots,M\right\}^{*}$, if $w_{n}<M$ then we set $w^{+}:=w_{1}\ldots w_{n-1}(w_{n}+1)$; if $w_{n}>0$ then we set $w^{-}:=w_{1}\ldots w_{n-1}(w_{n}-1)$. We also write $w^{\infty}$ for the periodic sequence with periodic block $w$. For a sequence $(c_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ we denote its reflection by $\overline{(c_{i})}=(M-c_{1})(M-c_{2})\ldots$. Furthermore, we use lexicographical order between sequences in $\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$. For two infinite sequences $(c_{i}),(d_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ we write $(c_{i})\prec(d_{i})$ if there exists $N\in\mathbb{N}$ such that $c_{1}\ldots c_{N-1}=d_{1}\ldots d_{N-1}$ and $c_{N}<d_{N}$. We also write $(c_{i})\preccurlyeq(d_{i})$ if $(c_{i})\prec(d_{i})$ or $(c_{i})=(d_{i})$. Similarly, we write $(c_{i})\succ(d_{i})$ if $(d_{i})\prec(c_{i})$; and write $(c_{i})\succcurlyeq(d_{i})$ if $(d_{i})\preccurlyeq(c_{i})$.
@@
$0^{\infty}\prec\sigma^{n}((\delta_{i}))\preccurlyeq(\delta_{i})\quad\forall n\geq 0$
@@
In general, for $\beta\in(1,M+1)$ and $x\in(0,1]$ let $a(x,\beta)$ be the quasi-greedy $\beta$-expansion of $x$, i.e., the lexicographically largest sequence $(a_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ not ending with $0^{\infty}$ and satisfying $\pi_{\beta}((a_{i}))=x$. The following result for quasi-greedy expansions was established by Parry [26] and de Vries, Komornik and Loreti [10, Lemma 2.2 and Proposition 2.4].
@@
$\left\{(a_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}:0^{\infty}\prec\sigma^{n}((a_{i}))\preccurlyeq\delta(\beta)\leavevmode\nobreak\ \forall n\geq 0\right\}.$
@@
For $\beta\in(1,M+1)$ we recall from (1.2) and (1.3) the definitions of intrinsic univoque set $\widetilde{U}_{\beta,M}$ and univoque set ${U}_{\beta,M}$, respectively. According to the coding map $\Pi_{\beta}$ in (1.1), we define the symbolic intrinsic univoque set $\widetilde{{\mathbf{U}}}_{\beta,M}\subset{\Omega}_{M}^{\mathbb{N}}$ and symbolic univoque set ${\mathbf{U}}_{\beta,M}\subset{\Omega}_{M}^{\mathbb{N}}$ by
@@
The following characterization of $\widetilde{{\mathbf{U}}}_{\beta,M}$ plays an important role in the proof of Theorem 1.4. When $M=1$, it was proven in [23, Proposition 2.2].
@@
In view of Figure 1, there are four possible patterns for $f_{\beta,d}(\Delta_{\beta,M})\setminus O_{\beta,M}$ with $d\in\Omega_{M}$: see Figure 3 for the typical pattern, and see Figure LABEL:fig:1' for the patterns in the three corners. Since $d_{n}^{1}=s<M$, this implies $\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)<\frac{s+1}{\beta}$. Note that $d_{n}^{1}=s$. Then
@@
Suppose on the contrary that $d_{n+1}^{1}d_{n+2}^{1}\ldots\succcurlyeq\delta(\beta)$. If the equality holds, then we have $\pi_{\beta}(d_{n+1}^{1}d_{n+2}^{1}\ldots)=1$, leading to a contradiction with (2.4). If strict inequality holds, then there exists $k\in\mathbb{N}$ such that $d_{n+1}^{1}\ldots d_{n+k-1}^{1}=\delta_{1}(\beta)\ldots\delta_{k-1}(\beta)$ and $d_{n+k}^{1}>\delta_{k}(\beta)$. By Lemmas 2.1 and 2.2 it follows
@@
again leading to a contradiction with (2.4). This proves $d_{n+1}^{1}d_{n+2}^{1}\ldots\prec\delta(\beta)$ as required.
@@
Note that $d_{n}^{1}=s<M$. Starting with $k_{0}=n$ we define by (2.3) an infinite sequence of indices $k_{0}<k_{1}<k_{2}<\cdots$ satisfying for $j=1,2,\ldots$ the conditions
$d_{k_{j-1}+1}^{1}\ldots d_{k_{j}-1}^{1}=\delta_{1}(\beta)\ldots\delta_{k_{j}-k_{j-1}-1}(\beta)\quad\textrm{and}\quad d_{k_{j}}^{1}<\delta_{k_{j}-k_{j-1}}(\beta).$
Then by Lemma 2.1 it follows that

### v2

# Phase transitions for unique codings of fat Sierpinski Gaskets

@@ Abstract.
Given an integer $M\geq 1$ and $\beta\in(1,M+1)$, let $S_{\beta,M}$ be the fat Sierpinski gasket in $\mathbb{R}^{2}$ generated by the iterated function system $\left\{f_{d}(x)=\frac{x+d}{\beta}:d\in\Omega_{M}\right\}$, where $\Omega_{M}=\{(i,j)\in\mathbb{Z}_{\geq 0}^{2}:i+j\leq M\}$. Then each $x\in S_{\beta,M}$ may be represented as a series $x=\sum_{i=1}^{\infty}\frac{d_{i}}{\beta^{i}}=:\Pi_{\beta}((d_{i}))$, and the infinite sequence $(d_{i})\in\Omega_{M}^{\mathbb{N}}$ is called a coding of $x$. Since $\beta<M+1$, a point in $S_{\beta,M}$ may have multiple codings. Let $U_{\beta,M}$ be the set of $x\in S_{\beta,M}$ having a unique coding, that is
@@
When $M=1$, Kong and Li [2020, Nonlinearity] described two critical bases for the phase transitions of the intrinsic univoque set $\widetilde{U}_{\beta,1}$, which is a subset of $U_{\beta,1}$. In this paper we consider $M\geq 2$, and characterize the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ for the phase transitions of $U_{\beta,M}$: (i) if $\beta\in(1,\beta_{G}(M)]$, then $U_{\beta,M}$ is finite; (ii) if $\beta\in(\beta_{G}(M),\beta_{c}(M))$ then $U_{\beta,M}$ is countably infinite; (iii) if $\beta=\beta_{c}(M)$ then $U_{\beta,M}$ is uncountable and has zero Hausdorff dimension; (iv) if $\beta>\beta_{c}(M)$ then $U_{\beta,M}$ has positive Hausdorff dimension. Moreover, we show that the first critical base $\beta_{G}(M)$ is a Perron number, while the second critical base $\beta_{c}(M)$ is a transcendental number.
@@
Fat Sierpinski gasket, unique coding, phase transition, critical base, Thue-Morse sequence.
@@ 1. Introduction
Note that each digit $\alpha_{ij}=(i,j)$ is a vector in $\mathbb{R}^{2}$ which determines the transition of $f_{\beta,\alpha_{ij}}$. It is well-known that $S_{\beta,M}$ is the unique nonempty compact set satisfying the equation $S_{\beta,M}=\bigcup_{d\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega_{M}}}f_{\beta,d}(S_{\beta,M})$ (cf. [18]). To study $S_{\beta,M}$ it is useful to introduce the coding map
@@
which is always continuous and surjective. If $\beta>M+1$, the IFS $\mathscr{F}_{\beta,M}=\left\{f_{\beta,d}:d\in\Omega_{M}\right\}$ satisfies the strong separation condition, and then $\Pi_{\beta}$ is bijective (cf. [15]). If $\beta=M+1$, the IFS $\mathscr{F}_{\beta,M}$ fails the strong separation condition but still admits the open set condition, and thus $\Pi_{\beta}$ is nearly bijective up to a countable set. However, for $\beta<M+1$ the IFS $\mathscr{F}_{\beta,M}$ fails the open set condition and it always has overlaps. In particular, if $\beta\in(1,\frac{M+2}{2}]$ the attractor $S_{\beta,M}$ is an isosceles triangle with vertices $(0,0),(\frac{M}{\beta-1},0)$ and $(0,\frac{M}{\beta-1})$ (see Proposition 2.5). In this case, for Lebesgue almost every $x\in S_{\beta,M}$ the preimage $\Pi_{\beta}^{-1}(x)$ is uncountable (cf. [29]). When $\beta\in(\frac{M+2}{2},M+1)$ the structure of $S_{\beta,M}$ is getting more complicated. Broomhead et al. [5] showed that for $M=1$ and $\beta\leq\beta_{*}$ the set $S_{\beta,M}$ has non-empty interior, where $\beta_{*}\approx 1.54369$ is a root of $x^{3}-2x^{2}+2x-2=0$.
@@
[…]
@@
(1.2) $\widetilde{U}_{\beta,M}:=\left\{\Pi_{\beta}((d_{i}))\in S_{\beta,M}:\Pi_{\beta}(d_{n+1}d_{n+2}\ldots)\notin O_{\beta,M}\penalty 10000\ \forall n\geq 0\right\},$
where $\Pi_{\beta}$ is defined in (1.1). Note that $\widetilde{U}_{\beta,M}$ is the survivor set of some open dynamical systems (cf. [28]). Indeed, let $T_{\beta,M}$ be the expanding map from the isosceles triangle $\Delta_{\beta,M}$ to itself such that $T_{\beta,M}(x)=f_{\beta,d}^{-1}(x)$ if $x\in f_{\beta,d}(\Delta_{\beta,M})$ for any $d\in{\Omega}_{M}$. Then $S_{\beta,M}$ is invariant under $T_{\beta,M}$, and the intrinsic univoque set $\widetilde{U}_{\beta,M}$ is the survivor set consisting of all $x\in\Delta_{\beta,M}$ whose orbit under $T_{\beta,M}$ never enters the hole $O_{\beta,M}$, i.e., $\widetilde{U}_{\beta,M}=\left\{x\in\Delta_{\beta,M}:T_{\beta,M}^{n}(x)\notin O_{\beta,M}\penalty 10000\ \forall n\geq 0\right\}$. For more information on open dynamical systems we refer to the papers [31, 12, 13, 11, 16, 7, 6, 20, 1] and the references therein.
@@
where #A\#A denotes the cardinality of a set $A$. Note that each point in $\widetilde{U}_{\beta,M}$ has a unique coding under $\Pi_{\beta}$. Then $\widetilde{U}_{\beta,M}\subset U_{\beta,M}$ for any $\beta\in(1,M+1)$.
@@
Write $\mathbb{N}:=\left\{1,2,3,\ldots\right\}$, $\mathbb{N}_{0}:=\mathbb{N}\cup\left\{0\right\}$. Let $(\tau_{i})_{i=0}^{\infty}=01101001\ldots\in\left\{0,1\right\}^{\mathbb{N}}$ be the classical Thue-Morse sequence defined by (cf. [2])
@@
where $(\tau_{i})_{i=0}^{\infty}=01101001\ldots$ is the classical Thue-Morse sequence. Similarly, if $M=3N+2$ with $N\in\mathbb{N}_{0}$, then $\beta_{c}(3N+2)$ is the unique root in $(N+1,N+2)$ of the equation
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[\frac{M}{3}\right]}+1\leq\beta_{G}(M)<\beta_{c}(M)<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left[\frac{M}{3}\right]}+2,$
@@
We will show that each $\beta_{G}(M)$ is a Perron number, and each $\beta_{c}(M)$ is a transcendental number. When $M=1$ the following phase transition phenomenon for $\widetilde{U}_{\beta,1}$ was established by Sidorov [29, Theorem 4.1] for the cardinality transition, and by Kong and Li [24, Theorem 2] for the dimension transition.
@@
    If $\beta\in(1,\beta_{G}(1)]$, then $\widetilde{U}_{\beta,1}=\#U_{\beta,1}=3$;
@@
In this paper we extend Theorem 1.3 to $M\geq 2$, and show that (i) $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$ are finite if and only if $\beta\leq\beta_{G}(M)$; (ii) $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$ have positive Hausdorff dimension if and only if $\beta>\beta_{c}(M)$. The new challenge is that for $M=3N+1$ with $N\geq 1$ the digit N1N-1 might occur in $d_{n}^{1}$ or $d_{n}^{2}$ for a sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$, while for $M=1$ this can never happen. Fortunately, we are able to show that this digit N1N-1 can occur at most 3 times in any sequence of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ (see Lemma 4.5). Furthermore, completely different from the case for $M=1$, we need more effort to find the critical base $\beta_{c}(M)$ for $M=3N+3$. In other words, we need to construct a completely new type of sequences to represent $\beta_{c}(M)$; see more in Section 6.
@@
    If $\beta\in(1,\beta_{G}(M)]$, then $U_{\beta,M}=\#\widetilde{U}_{\beta,M}=3$;
@@
    If $\beta>\beta_{c}(M)$, then both $U_{\beta,M}$ and $\widetilde{U}_{\beta,M}$ have positive Hausdorff dimension.
@@
The rest of the paper is organized as follows. In the next section we give the characterization of $\widetilde{U}_{\beta,M}$. In Section 3 we show that the generalized golden ratio $\beta_{G}(M)$ defined in Definition 1.1 is the first critical base for $\widetilde{U}_{\beta,M}$ in the sense that $\widetilde{U}_{\beta,M}$ is finite if and only if $\beta\leq\beta_{G}(M)$. Furthermore, we show that the generalized Komornik-Loreti constant $\beta_{c}(M)$ in Definition 1.2 is the second critical base for $\widetilde{U}_{\beta,M}$. This will be proved in Sections 46 for $M=3N+1,M=3N+2$ and $M=3N+3$, respectively. We point out that the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ are also applied to $U_{\beta,M}$ by Proposition 2.5.
@@ 2. Characterization of $\widetilde{U}_{\beta,M}$
This representation plays an important role in our proof of Theorem 1.4; and it will be clear after we prove Proposition 2.3. Note by (2.1) that each row in the above presentation is a length $n$ word or an infinite sequence over the alphabet $\left\{0,1,\ldots,M\right\}$.
@@
Note that the two induced sequences $(d_{i}^{1})$ and $(d_{i}^{2})$ both have digits in $\left\{0,1,\ldots,M\right\}$. Then we also need to introduce the one-dimensional symbolic dynamics $(\left\{0,1,\ldots,M\right\}^{\mathbb{N}},\sigma)$, where $\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ is the set of all infinite sequences with each digit from $\left\{0,1,\ldots,M\right\}$, and $\sigma$ is the left-shift map. For $n\in\mathbb{N}$ let $\left\{0,1,\ldots,M\right\}^{n}$ be the set of all length $n$ words with each digit from $\left\{0,1,\ldots,M\right\}$. Set $\left\{0,1,\ldots,M\right\}^{*}:=\bigcup_{n=1}^{\infty}\left\{0,1,\ldots,M\right\}^{n}$. For a word $w=w_{1}\ldots w_{n}\in\left\{0,1,\ldots,M\right\}^{*}$, if $w_{n}<M$ then we set $w^{+}:=w_{1}\ldots w_{n-1}(w_{n}+1)$; if $w_{n}>0$ then we set $w^{-}:=w_{1}\ldots w_{n-1}(w_{n}-1)$. We also write $w^{\infty}$ for the periodic sequence with the periodic block $w$. For a sequence $(c_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ we denote its reflection by $\overline{(c_{i})}=(M-c_{1})(M-c_{2})\ldots$. Furthermore, we use lexicographical order between sequences in $\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$. For two infinite sequences $(c_{i}),(d_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ we write $(c_{i})\prec(d_{i})$ if there exists $N\in\mathbb{N}$ such that $c_{1}\ldots c_{N-1}=d_{1}\ldots d_{N-1}$ and $c_{N}<d_{N}$. We also write $(c_{i})\preccurlyeq(d_{i})$ if $(c_{i})\prec(d_{i})$ or $(c_{i})=(d_{i})$. Similarly, we write $(c_{i})\succ(d_{i})$ if $(d_{i})\prec(c_{i})$; and write $(c_{i})\succcurlyeq(d_{i})$ if $(d_{i})\preccurlyeq(c_{i})$.
@@
$0^{\infty}\prec\sigma^{n}((\delta_{i}))\preccurlyeq(\delta_{i})\quad\forall n\geq 0.$
@@
In general, for $\beta\in(1,M+1)$ and $x\in(0,1]$ let $a(x,\beta)$ be the quasi-greedy $\beta$-expansion of $x$, i.e., the lexicographically largest sequence $(a_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ not ending with $0^{\infty}$ and satisfying $\pi_{\beta}((a_{i}))=x$. The following result for quasi-greedy expansions was established by Parry [27] and de Vries et al. [10, Lemma 2.2 and Proposition 2.4].
@@
$\left\{(a_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}:0^{\infty}\prec\sigma^{n}((a_{i}))\preccurlyeq\delta(\beta)\penalty 10000\ \forall n\geq 0\right\}.$
@@
For $\beta\in(1,M+1)$ we recall from (1.2) and (1.3) the definitions of the intrinsic univoque set $\widetilde{U}_{\beta,M}$ and univoque set ${U}_{\beta,M}$, respectively. According to the coding map $\Pi_{\beta}$ in (1.1), we define the symbolic intrinsic univoque set $\widetilde{{\mathbf{U}}}_{\beta,M}\subset{\Omega}_{M}^{\mathbb{N}}$ and the symbolic univoque set ${\mathbf{U}}_{\beta,M}\subset{\Omega}_{M}^{\mathbb{N}}$ by
@@
The following characterization of $\widetilde{{\mathbf{U}}}_{\beta,M}$ plays an important role in the proof of Theorem 1.4. When $M=1$, it was proven in [24, Proposition 2.2].
@@
Before proving the proposition we clarify our main idea. Note that $d_{n}^{1}<M$ implies that the first coordinate of $\Pi_{\beta}(d_{n}d_{n+1}\ldots)\in U_{\beta,M}$ is strictly smaller than $\frac{d_{n}^{1}+1}{\beta}$. Similar to the one-dimensional unique $\beta$-expansions we can deduce that this is equivalent to $d_{n+1}^{1}d_{n+2}^{1}\ldots\prec\delta(\beta)$. Similar ideas can be applied to study the case for $d_{n}^{2}<M$. However, the case for $\overline{d_{n}^{\oplus}}<M$ is more involved. Note that $\overline{d_{n}^{\oplus}}=M-(d_{n}^{1}+d_{n}^{2})$. Then $\overline{d_{n}^{\oplus}}<M$ is equivalent to $d_{n}^{1}+d_{n}^{2}>0$. Observe that $d_{n}\in\Omega_{M}$, i.e., $d_{n}^{1},d_{n}^{2}\in\left\{0,1,\ldots,M\right\}$ and $d_{n}^{1}+d_{n}^{2}\leq M$. Similar to the previous two cases, we split the proof of the case for $\overline{d_{n}^{\oplus}}<M$ into four subcases: (i) $d_{n}^{1}<M,d_{n}^{2}<M$ and $d_{n}^{1}+d_{n}^{2}>0$; (ii) $d_{n}^{1}+d_{n}^{2}=0$, i.e., $d_{n}^{1}=d_{n}^{2}=0$; (iii) $d_{n}^{1}=0$ and $d_{n}^{2}=M$; (iv) $d_{n}^{1}=M$ and $d_{n}^{2}=0$. Each subcase can be proved similar to the one-dimensional unique $\beta$-expansions.
@@
In view of Figure 1, there are four possible patterns for $f_{\beta,d}(\Delta_{\beta,M})\setminus O_{\beta,M}$ with $d\in\Omega_{M}$: see Figure 3 for the typical pattern, and see Figure 4 for the patterns in the three corners. Since $d_{n}^{1}=s<M$, this implies $\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)<\frac{s+1}{\beta}$. Note that $d_{n}^{1}=s$. Then
@@
Suppose on the contrary that $d_{n+1}^{1}d_{n+2}^{1}\ldots\succcurlyeq\delta(\beta)$. If the equality holds, then we have $\pi_{\beta}(d_{n+1}^{1}d_{n+2}^{1}\ldots)=1$, leading to a contradiction to (2.4). If strict inequality holds, then there exists $k\in\mathbb{N}$ such that $d_{n+1}^{1}\ldots d_{n+k-1}^{1}=\delta_{1}(\beta)\ldots\delta_{k-1}(\beta)$ and $d_{n+k}^{1}>\delta_{k}(\beta)$. By Lemmas 2.1 and 2.2 it follows that
@@
again leading to a contradiction to (2.4). This proves $d_{n+1}^{1}d_{n+2}^{1}\ldots\prec\delta(\beta)$ as required.
@@
Note that $d_{n}^{1}=s<M$. Starting with $k_{0}=n$ we define by (2.3) an infinite sequence of indices $k_{0}<k_{1}<k_{2}<\cdots$ satisfying the conditions
$d_{k_{j-1}+1}^{1}\ldots d_{k_{j}-1}^{1}=\delta_{1}(\beta)\ldots\delta_{k_{j}-k_{j-1}-1}(\beta)\quad\textrm{and}\quad d_{k_{j}}^{1}<\delta_{k_{j}-k_{j-1}}(\beta)$
for $j=1,2,\ldots$. Then by Lemma 2.1 it follows that